OpenStudy (anonymous):

3.5x-20=2.4x+13 solve for x

OpenStudy (anonymous):

is (.) a multiplication

OpenStudy (anonymous):

no it's a decimal

OpenStudy (anonymous):

start by adding 20 to both sides then

OpenStudy (anonymous):

3.5x-2.4x=13+20 1.1x=33 x=33/1.1 x=30

OpenStudy (anonymous):

thank you:)